# Feather the Throttle: Revisiting Visual Token Pruning for Vision-Language Model Acceleration

### v1

@@ Abstract
Recent works on accelerating Vision-Language Models show that strong performance can be maintained across a variety of vision-language tasks despite highly compressing visual information. In this work, we examine the popular acceleration approach of early pruning of visual tokens inside the language model and find that its strong performance across many tasks is not due to an exceptional ability to compress visual information, but rather the benchmarks’ limited ability to assess fine-grained visual capabilities. Namely, we demonstrate a core issue with the acceleration approach where most tokens towards the top of the image are pruned away. Yet, this issue is only reflected in performance for a small subset of tasks such as localization. For the other evaluated tasks, strong performance is maintained with the flawed pruning strategy. Noting the limited visual capabilities of the studied acceleration technique, we propose FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria), a straightforward approach that (1) resolves the identified issue with early-layer pruning, (2) incorporates uniform sampling to ensure coverage across all image regions, and (3) applies pruning in two stages to allow the criteria to become more effective at a later layer while still achieving significant speedup through early-layer pruning. With comparable computational savings, we find that FEATHER has more than $\mathbf{5\times}$ performance improvement on the vision-centric localization benchmarks compared to the original acceleration approach.
@@ 1 Introduction
Figure 1: (a) Although FastV prunes most visual tokens from the upper portion of the image, the approach still displays strong performance on a variety of evaluated vision-language tasks except for the vision-centric task of localization. (b) Based on our findings, we propose FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria), a straightforward approach that resolves the existing issue of selecting bottom tokens, additionally maintains uniformly sampled tokens to ensure good coverage over the whole image, and prunes in two stages.
@@
The motivation of our study is to get a better understanding of the vision capabilities of accelerated VLMs given that they leverage highly compressed visual information, focusing specifically on FastV approach. When evaluating across a wide range of vision-language tasks, we find that while the approach maintains strong performance across many tasks, compression causes a substantial decrease in performance for TextVQA and a severe performance for localization tasks. While this result is not particularly surprising as visual grounding is expected to suffer when compressing visual information, when we analyze the approach’s poor performance we uncover a fundamental issue with the approach that the pruning criteria is ineffective in early layers, heavily discarding tokens towards the top of the image. As this issue is not specific to any particular task, we then study how this cost is hidden on the majority of tasks and find that most evaluated benchmarks require minimal visual grounding, even for many that show a substantial gap between vision enabled and disabled setups. This finding underscores a significant challenge in the field of multimodal learning, not only for measuring the effectiveness of VLM acceleration methods but also for benchmarking the visual capabilities of VLMs as a whole.
@@
Given the discovered limitation of the studied VLM acceleration approach, we next experiment with various alternative criteria and design choices to better compress visual information, leading to our final method, FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria). Specifically, we prune after an early layer with a modified version of the criteria that resolves the uneven selection issue and also maintain a small amount of uniformly sampled tokens to ensure adequate coverage over all image regions. Furthermore, we apply more extensive pruning at a later layer, where the effectiveness of the attention-based criteria is enhanced. This strategy is analogous to how a racecar driver feathers the throttle by gradually pressing the accelerator at the beginning of a turn to maintain grip and then accelerating more aggressively once the car is past the apex. We show that FEATHER results in substantial performance gains compared to the original acceleration approach, improving localization performance more than $\mathbf{5\times}$ with comparable computational savings. Strikingly, we find that our approach achieves this performance improvement while only retaining 3.3% of visual tokens for the second half of LLM layers. Overall, our work demonstrates that while visual compression can maintain strong performance even on challenging vision-centric tasks, its effectiveness depends on a well-designed strategy, which is currently difficult to assess due to many vision-language benchmarks not thoroughly evaluating vision capabilities.
@@ 2 Related Work
Recent efforts to accelerate VLMs can be broadly divided into two main categories: compressing visual information before it enters the LLM, and compressing visual information within the LLM itself. In the first category, Chat-UniVi [13] dynamically merges visual tokens with similar semantic meanings. Alternatively, PruMerge [25] selects important tokens according to the similarities between the class token from the vision encoder and the individual patch tokens and then merges important tokens with the remaining unselected tokens by a weighted average. For the LLaVA-NeXT [23] approach of partitioning an image into sub-images where inefficiency is an even bigger problem, HiRED [3] selects tokens with top feature important on each sub-image with an allocated budget. Other methods argue that the input image alone does not include enough information to select important patches and thus use the textual input to recover visually meaningful tokens [8, 34].
@@
3 The Drawback of VLM Acceleration
@@ 3.1 Preliminaries
Open-Ended Visual Question Answering. We evaluate general visual reasoning using the VizWiz [5] and VQAv2 [9] datasets, spatial reasoning using the GQA [12] dataset, and reasoning around text using the TextVQA [26] dataset.
@@
Figure 2: Contrasting the difference in performance dropoff on the challenging vision-centric localization task (Left) versus the other evaluated tasks (Middle) when pruning visual tokens after the shallow LLM layers. Whereas performance decrease is minimal for most tasks, localization exhibits roughly a linear decrease to zero as the ratio of pruned tokens increases. Right: Per-task performance breakdown across various setups of pruning ratios. Using $\textit{K}=3$ for all pruning setups.
3.2 Early visual token pruning falters in vision-centric tasks
@@
We find that for the majority of evaluated tasks, heavily reducing the number of visual tokens after the shallow LLM layers results in minimal performance dropoff. Specifically, comparing the baseline model to the setup dropping 75% of tokens, we observe the following decreases in performance: 3.5% for AI2D, 0.1% for VSR, 3.7% for TallyQA, 4.8% for POPE, 5.1% for VizWiz, 7.9% for VQAv2 and 7.7% for GQA. By contrast, TextVQA shows a much more substantial drop, with a 42.0% decrease. Notably, the task with the highest performance dropoff is decidedly localization, with RefCOCO, RefCOCO+, RefCOCOg, and OCID-Ref exhibiting performance decreases of 88.9%, 88.9%, 91.0%, and 86.0%, respectively. Investigating further, Figure 2 shows that localization performance declines roughly linearly to zero as the pruning ratio progresses from 0 to 1, a pattern that sharply contrasts with most other tasks, where performance remains largely unaffected.
It is important to note that localization is undoubtedly a challenging task for a model that discards visual tokens. The task inherently requires visual grounding, as the model must not only identify the object of interest but also retain sufficient tokens corresponding to the object to accurately define its boundaries. Nonetheless, it is still surprising that (1) localization performance declines to the extent it does even when many tokens are still maintained (e.g., decreasing average performance by 45% when dropping 50% of tokens), and (2) most non-localization benchmarks do not exhibit a similar pattern whatsoever (except for TextVQA), despite these benchmarks aiming to evaluate visual grounding abilities, such as counting for TallyQA, spatial reasoning for GQA, and chart understanding for AI2D. In the following sections, we further explore these unexpected findings.
Finding 1: Pruning visual tokens after shallow LLM layers results in a jarring performance decline for localization benchmarks and a moderate decrease for TextVQA, whereas performance remains relatively unchanged for other evaluated tasks.
@@ 3.3 Interpreting poor vision-centric task performance
We next investigate why pruning visual tokens after the shallow LLM layers has such an adverse effect on localization performance. Notably, performance begins to decrease even with minimal token pruning, suggesting the pruning criteria, intended to retain important tokens, seems to not effectively capture the necessary visual information.
@@
To assess the functionality of the pruning criteria, we first examine the distribution of retained tokens across all benchmark examples (both localization and non-localization). As shown in Figure 3(b), we find that pruning visual tokens after the shallow LLM layers ($K=3$) retains tokens concentrated at the bottom of the image. In addition to evaluating the effectiveness of the criteria applied after the shallow LLM layers, we also examine the criteria behavior when applied after later layers ($K\in(8,16,24)$). Notably, we observe that as tokens are pruned later in the model, the criteria shift away from the bias of maintaining tokens based on visual token position. In addition, qualitatively we observe that as the pruning layer increases, not only is the positional bias reduced, but the criteria also begins to correctly select tokens relevant to the text instruction. For example, as shown in Figure 3(a), pruning after the later layers results in a distinct concentration of selected tokens around the area corresponding to the reference expression for localization.
To quantify this change in criteria effectiveness across layers, we measure the performance when varying the pruning layer ($K\in\{3,8,16,24\}$). As shown in Figure 3(c), performance remains stable when pruning is performed at deeper layers ($K=16$ and $K=24$), whereas pruning at earlier layers ($K=3$ and $K=8$) leads to performance degradation, particularly for localization. This result demonstrates that the current criteria based on attention received from the last instruction token can discern important visual tokens; however, it currently shows limitations that need to be addressed.
Finding 2: The pruning criteria when applied after shallow layers predominantly selects visual tokens from the bottom part of the image.
3.4 Explaining VLM inference acceleration performance on other tasks
Given that early pruning of visual tokens leads to a suboptimal tokens selection where remaining tokens are concentrated towards the bottom of the image, a natural question arises: how does the approach still maintain high performance across a diverse range of tasks, including those aiming to evaluate visual grounding abilities? In this section, we test several hypotheses explaining this puzzling observation, ultimately concluding that strong performance is caused by the limitation of the tasks to assess vision capabilities.
Consider the example presented at the top of Figure 1, where the question is, “what animal is in front of the trees?” In this case, even when all visual tokens corresponding to trees are pruned, the model still predicts the correct answer. One possible explanation is that, because pruning occurs within the LLM, the visual information from the pruned tokens is transferred to the retained tokens prior to pruning. This would also account for poor localization performance, as positional information, unlike visual content, cannot be transferred through the attention mechanism (using RoPE). Alternatively, the model may predict the correct answer because the answer can be inferred without direct visual grounding (in this example, despite all tree tokens being pruned, the model retains tokens for a cow).
@@ 3.4 Explaining VLM inference acceleration performance on other tasks
$K$ Criteria FLOPS Red OCID-Ref RefCOCOg RefCOCO+ RefCOCO Avg Other Task Avg
3 Attention-based
$\phi_{\text{original}}$ 68% 0.057 0.051 0.061 0.067 0.059 0.594
$\phi_{\textbf{{-R}}}$ 68% 0.229 0.151 0.133 0.153 0.167 0.618

### v2

@@ Abstract
Recent works on accelerating Vision-Language Models achieve strong performance across a variety of vision-language tasks despite highly compressing visual information. In this work, we examine the popular acceleration approach of early pruning of visual tokens inside the language model. Surprisingly, we find that while strong performance is maintained across many tasks, it exhibits drastically different behavior for a subset of vision-centric tasks such as localization. Upon further investigation, we uncover a core issue with the acceleration approach where most tokens towards the top of the image are pruned away. Yet, on many benchmarks aiming to evaluate vision-centric capabilities, strong performance persists with the flawed pruning strategy, highlighting these benchmarks’ limited ability to assess fine-grained visual capabilities. Based on these findings, we propose FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria), a straightforward approach that resolves the discovered early-layer pruning issue and further enhances the preservation of relevant tokens via multistage pruning with early uniform sampling to ensure broad image coverage. With comparable computational savings, we find that FEATHER achieves more than $\mathbf{5\times}$ performance improvement on the vision-centric localization benchmarks compared to the original acceleration approach.
@@ 1 Introduction
Figure 1: (a) Although FastV prunes most visual tokens from the upper portion of the image, the approach still displays strong performance on a variety of evaluated vision-language tasks except for a small subset of vision-centric tasks like localization. (b) Based on our findings, we propose FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria), a straightforward approach that resolves the existing issue of selecting bottom tokens, additionally maintains uniformly sampled tokens to ensure good coverage over the whole image, and prunes in two stages.
@@
The motivation of our study is to get a better understanding of the vision capabilities of accelerated VLMs given that they leverage highly compressed visual information, focusing specifically on FastV approach. Evaluating across a wide range of vision-language tasks, we find that while the approach maintains strong performance across many tasks, there is a substantial decrease in performance for TextVQA and a severe decrease for localization tasks. When analyzing the approach’s poor performance, we uncover a fundamental issue with the pruning criteria in early layers where tokens towards the top of the image are disproportionally removed due to the long-term decay property of RoPE. Since this issue is not task-specific, we examine how performance is maintained across most evaluated tasks and find that even when tokens are entirely discarded using the flawed criteria, performance remains largely unchanged. This highlights how many benchmarks require minimal visual grounding. Our finding underscores a significant challenge in the field of multimodal learning, not only for measuring the effectiveness of VLM acceleration methods but also for benchmarking the visual capabilities of VLMs as a whole.
@@
Given the discovered limitation of the studied approach, we next explore whether the attention-based criteria can be modified to enable robust visual token pruning, leading to our final method, FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria). Specifically, we propose a straightforward modification to the attention-based criteria by removing RoPE, demonstrating that this resolves the positional bias issue and greatly improves performance. Additionally, we incorporate uniform sampling in early token pruning to ensure adequate coverage of all image regions and apply more extensive pruning at a later layer, when the attention-based criteria better discerns important tokens. This strategy is analogous to how a racecar driver feathers the throttle by gradually pressing the accelerator at the beginning of a turn to maintain grip and then accelerating more aggressively once the car is past the apex. We show that FEATHER results in substantial performance gains compared to the original acceleration approach, improving localization performance more than $\mathbf{5\times}$ with comparable computational savings. Strikingly, we find that our approach achieves this performance improvement while only retaining 3.3% of visual tokens for the second half of LLM layers. Overall, our work demonstrates that while visual compression can maintain strong performance even on challenging vision-centric tasks, its effectiveness depends on a well-designed strategy, which is currently difficult to assess due to many vision-language benchmarks not thoroughly evaluating vision capabilities.
In summary, our contributions include:
  • Revealing inconsistent effect of early pruning: Our evaluations across 12 benchmarks show that early visual token pruning minimally affects performance on most tasks, but causes a substantial dropoff for TextVQA and a drastic decline for localization tasks.
  • Analyzing poor performance: We uncover that the attention-based criteria for selecting important tokens exhibits a bias towards bottom image tokens in early layers due to the long-term decay property of RoPE.
  • Identifying benchmark limitations in assessing token pruning: We find that performance remains strong even with suboptimal pruning criteria and removal of pruned tokens before transferring information in the LLM, demonstrating how many benchmarks fail to assess fine-grained visual capabilities.
  • Improving attention-based criteria: With the straightforward modification of removing RoPE, we demonstrate that attention can be used as an effective criteria for selecting tokens and performance can be further improved with simple strategies like incorporating uniform sampling and pruning more extensively at a later layer, when the criteria better discerns important tokens.
@@ 2 Related Work
Recent efforts to accelerate VLMs can be broadly divided into two main categories: compressing visual information before it enters the LLM, and compressing visual information within the LLM itself. In the first category, Chat-UniVi [17] dynamically merges visual tokens with similar semantic meanings. PruMerge [31], FasterVLM [44], and VisionZip [40] select important tokens based on ViT attention ([CLS] or average attention received). PruMerge then merges important tokens with unselected ones while VisionZip merges unselected tokens together. For the LLaVA-NeXT [28] approach of partitioning an image into sub-images where inefficiency is an even bigger problem, HiRED [3] selects tokens with top feature important on each sub-image with an allocated budget. Other methods argue that the input image alone does not include enough information to select important patches and thus use the textual input to recover visually meaningful tokens [11, 41]. Alternatively, some approaches reduce visual information within the ViT by merging or pruning tokens [8, 30, 6].
@@
3 Revising Visual Token Pruning
@@ 3.1 Preliminaries
Open-Ended Visual Question Answering. We evaluate general visual reasoning using the VizWiz [5] and VQAv2 [13] datasets, spatial reasoning using the GQA [16] dataset, and reasoning around text using the TextVQA [32] dataset. For TextVQA, following [18], we exclude OCR-system parsed input tokens to better evaluate the effect of pruning on visual capabilities.
@@
3.2 Impact of early visual token pruning varies drastically by task
@@
For the majority of evaluated tasks, heavily reducing the number of visual tokens results in minimal performance decrease (3.5% for AI2D, 0.1% for VSR, 3.7% for TallyQA, 4.8% for POPE, 5.1% for VizWiz, 7.9% for VQAv2 and 7.7% for GQA when dropping 75% of tokens). However, for a small subset of tasks–particularly TextVQA, and more severely, localization–performance drops substantially, with decreases of 42.0%, 88.9%, 88.9%, 91.0%, and 86.0% for TextVQA, RefCOCO, RefCOCO+, RefCOCOg, and OCID-Ref, respectively. Investigating further, Figure 2 shows that localization performance declines roughly linearly to zero as the pruning ratio progresses from 0 to 1, a pattern that strongly contrasts with most other tasks, where performance remains largely unaffected. While prior works report a much smaller decline in TextVQA performance [39, 9], they incorporate OCR-parsed input tokens into prompts, reducing the reliance on visual information.
Discussion. Although localization is undoubtedly challenging for a model that discards visual tokens, as it requires fine-grained visual grounding and precise boundary identification, it is still surprising that (1) localization performance declines to the extent it does even when many tokens are maintained (e.g., dropping 50% of tokens reduces average performance by 45%), and (2) most non-localization benchmarks, except for TextVQA, exhibit a drastically different pattern despite these benchmarks aiming to evaluate visual grounding abilities, such as counting for TallyQA, spatial reasoning for GQA, and chart understanding for AI2D. In the following sections, we seek to explain these unexpected findings and understand their implications.
@@ 3.3 Interpreting poor vision-centric task performance
We next investigate why pruning visual tokens after the shallow LLM layers has such an adverse effect on localization performance. Intuitively, even localization should not require all image tokens, as there are background tokens irrelevant to identifying specific objects that can be pruned. Thus, it is surprising that performance begins to degrade even with minimal token pruning. This suggests that the pruning criteria, intended to retain important tokens, fails to effectively distinguish which tokens contain necessary visual information.
@@
To assess the efficacy of the pruning criteria, we first examine the distribution of retained tokens across all benchmark examples. As shown in Figure 3(b), we find that pruning visual tokens after the shallow LLM layers ($K=3$) retains tokens concentrated at the bottom of the image. For example, when $R=0.75$, the average sampled token position across all datasets is situated 80.7% of the way down the image. A Chi-Square test confirms the non-uniformity of selected tokens’ y-positions ($p$-value $<$ 0.05). We also examine the criteria behavior when pruning after later layers ($K\in(8,16,24)$) and observe that the bias of selecting bottom image tokens is reduced.
In addition, we observe that as the pruning layer increases, not only is the positional bias reduced, but the criteria also begins to correctly select tokens relevant to the text instruction. For example, as shown in Figure 3(a), pruning after the later layers results in a distinct concentration of selected tokens around the area corresponding to the reference expression for localization. To quantify this change in criteria effectiveness across layers, we measure the performance when varying the pruning layer ($K\in\{3,8,16,24\}$). As shown in Figure 3(c), performance remains stable when pruning is performed at deeper layers ($K=16$ and $K=24$), whereas pruning at earlier layers ($K=3$ and $K=8$) leads to performance degradation, particularly for localization.
Discussion. While previous research attributes the poor performance of early token pruning to the low redundancy of image tokens in shallow layers [39], we reveal a critical insight: the variable performance across pruning layers is linked to the effectiveness of selecting important tokens, as early token pruning leads to suboptimal token selection biased toward bottom tokens. To understand why the criteria disproportionately retains bottom image tokens, we must consider the LM architecture. Specifically, token selection is based on the attention received from the last text token. Since the LM uses RoPE to encode positional information [33], it exhibits a long-term decay property. While this property is well-suited for language modeling, we uncover that it poses an issue for selecting visual tokens, where tokens appearing later in the flattened, raster-scan order have higher attention scores purely due to their position. In language modeling, prior work finds that there is greater emphasis on shorter-distance information in shallow layers [15]. We predict that the visually-conditioned LM exhibits the same behavior, explaining why the positional bias is most pronounced when pruning after shallow layers.
@@
3.4 Explaining strong performance on other tasks
Given that early pruning of visual tokens leads to a suboptimal tokens selection where remaining tokens are concentrated towards the bottom of the image, a natural question arises: how does the approach still maintain high performance across a diverse range of tasks, including those aiming to evaluate visual grounding abilities? In this section, we test the following contrasting explanations:
  1. 1.
    Important visual information from top image tokens is transferred to later tokens before pruning with the unidirectional attention of the LLM.
  2. 2.
    Many questions can still be inferred with access to only the suboptimal set of visual tokens primarily from the bottom of the image.
To evaluate these hypotheses across tasks, we save the tokens retained by the studied pruning approach and, in a subsequent run, remove all other tokens before they reach the LLM, ensuring that no visual information from the pruned tokens is retained. As an additional comparison, we also evaluate a text-only model in which all visual tokens are removed before entering the LLM. As shown in Figure 4, we find that across all evaluated tasks, removing pruned tokens before the LLM results in little to no performance decrease compared to the method that allows information transfer in shallow layers prior to pruning.
Discussion. This finding indicates that the high performance of early visual token pruning on many tasks does not stem from the effectiveness of visual information transfer in early layers but rather signifies that many benchmarks do not demand a detailed understanding of visual information to answer questions accurately. Note that for the majority of these benchmarks (except VizWiz and AI2D), performance substantially declines when not including any visual information, signifying that simply comparing vision-enabled and disabled setups is insufficient to determine a benchmark’s ability to assess visual grounding.
4 Improving Visual Token Pruning
Based on our findings from §3, we now seek to answer: can attention serve as an effective criteria for token pruning even in early layers? To this end, we propose a straightforward solution to removing the positional bias of the attention-based criteria and demonstrate that our modifications enable effective pruning even when applied after early LLM layers (§4.1). Guided by our insights, we present our final approach, FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria) (§4.2), and showcase its impressive visual capabilities while offering high computational efficiency (§4.3).
4.1 Assessing potential of attention-based criteria
An attention-based criteria offers the potential to naturally identify and retain only the most relevant visual tokens for answering a question. However, as shown in §3, positional bias strongly influences token selection, limiting its effectiveness. Thus, we aim to resolve the positional bias issue and explore the true promise of attention-based criteria. We compare against non-attention-based approaches that incorporate an explicit inductive bias toward reducing redundancy. Below, we outline the explored criteria.
$\phi_{\textbf{{-R}}}$: Given the long-term decay property of RoPE, we propose the following straightforward adjustment to the pruning criteria in order to mitigate this bias and more effectively select tokens relevant to the text instruction. Namely, we still compute the attention score received from the last text token, except we do not apply RoPE to the attention mechanism, thereby removing the long-term decay effect. Note that this lightweight calculation is only done once per pruning, and all other state computations are compatible with FlashAttention [12].
$\phi_{\text{KNN}}$: Inspired by dynamic visual tokenization [17], we select tokens based on local density. For each visual token $z_{img}[i]$ where $i\in\{0,1,...,n-1\}$, we compute its local density $\rho_{i}$ using K-nearest neighbors and calculate distance index $\delta_{i}$ as:
[…]
$\delta_{i}$ represents how far away $z_{img}[i]$ is from other high-density tokens. We use $\rho_{i}*\delta_{i}$ as the token importance score for the criteria and select 196 tokens.
$\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$: Since $\phi_{\textbf{{-R}}}$ focuses on selecting important tokens and $\phi_{\text{uniform}}$ aims to reduce redundancy, we propose an ensemble of these two to leverage their respective strengths. Specifically, we apply $\phi_{\textbf{{-R}}}$ while incorporating a small number of uniformly sampled tokens using a stride of three.
Improvement of removing RoPE: In Table 1, we report results evaluating the effectiveness of these criteria. Comparing $\phi_{\textbf{{-R}}}$ with $\phi_{\text{original}}$, we find that by removing RoPE, at $K=3$ there is a 183% average improvement on localization tasks and at $K=8$, there is a 17% average improvement. These performance improvements demonstrate that once the impact of token position on attention is removed, attention score can more effectively be used for the criteria when applied after early LLM layers. Note that the narrowing gap in performance between $\phi_{\textbf{{-R}}}$ and $\phi_{\text{original}}$ from $K=3$ to $K=8$ is likely because $\phi_{\text{original}}$ has less of a bias towards selecting bottom image tokens as the pruning layer increases.
Localization Open-Ended VQA Challenge Sets
Pruning Layer Criteria
FLOPS Red X
Avg
OCID-Ref
RefCOCOg
RefCOCO+
RefCOCO
Avg
TextVQA
GQA
VQAv2
VizWiz
Avg
POPE
TallyQA
VSR
AI2D
Attention-based
$K=3$ $\phi_{\text{original}}$ 68% 5.9 5.7 5.1 6.1 6.7 54.8 31.8 58.4 72.7 56.3 64.0 83.2 57.1 63.3 52.4
$\phi_{\textbf{{-R}}}$ (Ours) 68% 16.7 22.9 15.1 13.3 15.3 59.0 41.6 61.2 76.0 57.3 64.7 85.2 58.2 62.2 53.2
$\Delta$ +10.7 +17.2 +10.0 +7.3 +8.6 +4.2 +9.7 +2.8 +3.2 +1.1 +0.7 +1.9 +1.1 -1.1 +0.8